# First-Order Phase Transition of the Schwinger Model with a Quantum Computer

### 20 Dec 2023 v1

@@ I Introduction
Exploring the phase diagram of a theory is central for the understanding of the fundamental laws of physics in many disciplines ranging from condensed matter to particle physics. For instance, in the case of ferromagnetic materials, critical points and phase boundaries help to understand and design magnetic materials with various applications [1, 2]. Within superconductivity, phase diagrams allow for identifying the critical temperatures and magnetic fields required for materials to exhibit zero electrical resistance [3, 4]. In particle physics, the phase diagram of Quantum Chromodynamics (QCD) elucidates the behavior of matter under extreme conditions, relevant for the early universe and neutron stars [5]. In high-energy physics, topological terms play an important role in the study of interesting phenomena such as the breaking of charge conjugation-parity symmetry in QCD [6], and the occurrence of out-of-equilibrium dynamical effects involving axion fields [7].
In order to explore the phase structure of a given theory, one often has to resort to numerical methods. A standard tool is Monte Carlo (MC) simulation, which has been successful for probing phase diagrams [8, 9, 10, 11], including that of lattice QCD [12]. However, the sign problem poses a barrier in certain parameter regimes [13], leaving relevant questions unanswered. Specifically for lattice QCD, large baryon chemical potentials or the presence of a topological $\theta$-term would trigger this problem [14]. Thus, variations of the conventional MC approach have been proposed to tackle this obstacle [12], but so far only with limited success. Hence, there is great interest in alternative methods that bypass the issue, such as tensor networks and quantum computing [15, 16, 17]. In particular, quantum computing offers a promising alternative route [16, 17, 18], with already a number of successful demonstrations [19, 20, 21, 22, 23, 24].
@@
Here we address these questions by studying two fermion discretizations, staggered and Wilson fermions, in order to explore any potential advantages of one type of fermion formulation over the other. We use the variational quantum eigensolver (VQE) [35] and derive a protocol for mapping this problem on a quantum circuit. The VQE, introduced as an alternative to quantum phase estimation [36], aligns with the capabilities of current and near-term quantum devices. We test our VQE using noiseless classical simulations between 6 to 12 qubits, to identify the best possible setup regarding the ansatz and gates that would capture the relevant ground states most efficiently. After the optimal ansatz-gate combination and variational parameters are found, the ground states across the phase transition are prepared on IBM’s quantum devices. We deomstrate that using state-of-the-art error mitigation techniques—zero noise extrapolation [37], readout error mitigation [38], Pauli twirling [39] and dynamical decoupling [40]—allows for obtaining precise results from the quantum measurements. To understand the minimum system sizes required to extrapolate faithfully to the continuum limit with a quantum computer, we use matrix product states (MPS). We numerically simulate intermediate system sizes and perform the continuum extrapolation, which we compare to analytical results. Our study also shows universality for the considered observables, as both discretizations lead to the same continuum values.
@@ II.2.1 Wilson fermions
$W_{W}&=x\sum_{n=0}^{N-2}\left(X_{2n+2}X_{2n+3}+Y_{2n+2}Y_{2n+3}\right)\\ &+\left(\frac{m_{\text{lat}}}{g}\sqrt{x}+x\right)\sum_{n=0}^{N-1}\left(X_{2n+1}X_{2n+2}+Y_{2n+1}Y_{2n+2}\right)\\ &+\sum_{n=0}^{N-2}\left(l_{0}+\sum_{k=0}^{n}Q_{k}\right)^{2}.\end{aligned}$ (9)
@@ III.1 Parametric ansatz circuits for VQE
Figure 2: Decomposition of a generic SO(4) gate depending on the six parameters $\theta_{1},\dots,\theta_{6}$ (a) the $R_{XX+YY}(\theta)=\exp(-i\theta(XX+YY)/2)$ (b), into CNOT and Pauli rotation gates. Boxes acting on a single qubit correspond to Pauli rotation gates, $R_{P}(\alpha)=\exp{-i\alpha P/2}$ ) with $P\in\{X,Y,Z\}$. Single-qubit gates where the argument is omitted refer to rotations around an angle $\pi/2$, $R_{P}(\pi/2)$. The light blue boxes represent the parameterized gates which are $R(\alpha,\beta,\gamma)=R_{X}(\gamma)R_{Z}(\beta)R_{X}(\alpha)$ in (a) and $R_{Y}(\theta)$ in (b). Panel (c) and (d) illustrate one layer of the brick and ladder ansatz, respectively, both following a non-parametric part for preparing the initial state |$\ket{\psi_{in}}$ (yellow box). The first layer in the brick ansatz has a CNOT-depth of $4$ whereas in ladder it is $22n-$ 2, where $n$ is the number of qubits, and in both cases it increases by $4$ with each layer.
@@ IV Results
Having summarized the methods in Sec. III that we have used to explore the lattice Schwinger model, we examine the performance of the ansätze for the VQE using classical simulations. Furthermore, we demonstrate the first-order phase transition with quantum hardware using the ground states obtained from the classcial simulation of the VQE. Finally, we perform the extrapolation to the continuum with data obtained from an MPS calculation and estimate the resource required for a reliable continuum limit.
@@ IV.1 VQE results
Focusing on the case of staggered fermions and using the initial states and qubit mappings discussed in Sec. III.1, the ansätze with the $R_{XX+YY}$ gates generally returned higher fidelities compared to the ones with the $SO(4)$ gates. In the latter case, we observed the tendency that the final parameters obtained by the optimization routine at the end of the VQE were all close to 00, resulting in a state formed only from the initial non-parametric part (which prepares the |$\ket{10}^{\otimes N/2}$ state). In contrast, with $R_{XX+YY}$ gates both of the architectures performed comparably well and were able to achieve fidelities above $0.99$. Since the brick architecture has smaller CNOT-depth, we again focus on this one for the staggered discretization. Considering $m_{\text{lat}}/g=10$ first, the staggered discretization behaves qualitatively similar to the Wilson fermions, as Fig. 3 reveals. We again observe the characteristic discontinuities in the electric field density (second row of Fig. 3) and the particle number (fourth row of Fig. 3), indicating the first-order quantum phase transition. Comparing to the Wilson fermions, the transition for the staggered discretization occurs at larger $l_{0}$, despite the fact that for both discretiaztions we match the qubit number, which effectively means that for the staggered fermions we have a larger number of physical sites and, since the volume is fixed to $N/\sqrt{x}=30$, a finer lattice spacing than for the Wilson fermions. This effect might be caused by a larger additive mass renormalization compared to the Wilson case. Also for the staggered discretization, we see the same trend that with increasing system size it is shifted to smaller values of the $\theta$-angle (for details, see Appendix D). In particular, the data from exact diagonalization is in excellent agreement with the one from simulated VQE with one ansatz layer, thus showing that our ansatz circuit together with the qubit mapping described in Sec. III is appropriate for capturing the physics of the model in the regime of large masses.
@@ IV.2 Inference runs on quantum hardware
Figure 3: Electric field density $\langle L_{W,S}\rangle$ and particle number $\langle P_{W,S}\rangle$ against $l_{0}$ with data from quantum hardware (black crosses), as compared to the noiseless expectation values (red pluses) and exact diagonalization (blue circles) for Wilson and staggered fermions respectively. The staggered fermions in this case have the same number of qubits as the Wilson fermions. Hence, while the title of each column specifies the $N$ for Wilson fermions, for staggered it is taken to be double that value. The lattice mass for these data is set to $m_{\text{lat}}/g=10$, so that we are above the second order phase transition of Fig. 1, without having to account for the mass shift. Therefore, we can observe the first-order phase transition. Note that the error bars are much smaller than the y-scale and thus, are not visible.
@@ IV.3 Continuum extrapolations with MPS
Figure 6: Electric field density $\mathcal{F}/{g}$ against $l_{0}$ for Wilson fermions and staggered fermions at $m_{\text{r}}/g=0.01$. The following MSE values are the mean squared error of each of the data as compared to the continuum mass perturbation theory prediction for the electric field density $\mathcal{F}/g$ given in Eq. (20). For staggered fermions with $\text{MS}_{t}$ and same $x$ (number of qubits as Wilson), the MSE is $1.041\times 10^{-6}$ ($6.16\times 10^{-7}$), and respectively for $\text{MS}_{L}$ $6.143\times 10^{-9}$ ($2.617\times 10^{-8}$). For Wilson fermions the MSE is $7.926\times 10^{-9}$.
@@ V Conclusion
Given that for the staggered case our MPS simulations showed better convergence towards the continuum limit, and our ansatz circuit for the VQE is slightly simpler, this suggests that the staggered discretization might be easier to address on current and near-term quantum hardware. Combining our results with some of the techniques of Ref. [19], it seems possible to reach larger lattice sizes on the order of 100 qubits, and for the future we plan to study the continuum extrapolation directly on quantum hardware. Moreover, it is an interesting question how both fermion discretizations perform in higher dimensions. While for the Schwinger model we integrated the gauge fields out, this is no longer possible for two and more spatial dimensions. This might change the performance of both discretiatzion and the complexity of a VQE anatz required to reliably capture the ground state.
Acknowledgements.
This work is partly funded by the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie COFUND scheme with grant agreement no. 101034267. P.N. acknowledges financial support from the Cyprus Research and Innovation Foundation under projects “Future-proofing Scientific Applications for the Supercomputers of Tomorrow (FAST)”, contract no. COMPLEMENTARY/0916/0048, and “Quantum Computing for Lattice Gauge Theories (QC4LGT)”, contract no. EXCELLENCE/0421/0019. This work is funded by the European Union’s Horizon Europe Framework Programme (HORIZON) under the ERA Chair scheme with grant agreement no. 101087126. This work is supported with funds from the Ministry of Science, Research and Culture of the State of Brandenburg within the Centre for Quantum Technologies and Applications (CQTA). A. C. is supported in part by the Helmholtz Association —“Innopool Project Variational Quantum Computer Simulations (VQCS).” The authors also thank the QC4HEP Working Group for discussions. IBM, the IBM logo, and ibm.com are trademarks of International Business Machines Corp., registered in many jurisdictions worldwide. Other product and service names might be trademarks of IBM or other companies.
@@ Appendix A Derivation of the Hamiltonian
In this appendix we derive the Hamiltonian for Wilson fermions, in particular, we provide the details of the fermion ordering chosen. Subsequently, we show the Jordan Wingner transformation and the final Hamiltonians in spin formulation used in the simulations for both Wilson and staggered fermions.
@@
$+\left(\frac{m_{\text{lat}}}{g}\sqrt{x}+xr\right)\sum_{n=0}^{N-1}\left(X_{2n+1}X_{2n+2}+Y_{2n+1}Y_{2n+2}\right)$
@@ Appendix B Zero noise extrapolations
Figure 7: Electric field density $\langle L_{W,S}\rangle$ and particle number $\langle P_{W,S}\rangle$ against noise factor for zero noise extrapolations (ZNE) on the data (black crosses) obtained from quantum hardware for different parameters specified in the titles of each subplot with Wilson and staggered fermions. The red pluses are noiseless values of the observables and the blue circles are the exact diagonalization results. For the fits we fit a first-order polynomial to either all points (blue continuous line) or to points with noise factor 1, 3 (dashed orange line). We also fit a second order polynomial to all points (green dotted dashed line). The brown triangle is the weighted average of the extrapolated points of these 3 fits. The plots on the left $(a-d)$ are showing examples where the ZNE helped take the expectation values measured closer to the noiseless result, while on the right $(e-h)$ we show occasions where the ZNE might not be able to improve the result. However, overall Fig. 5 indicates that ZNE can significantly improve the results. Where the error bars are not visible, they are smaller than the markers and y-scale.
@@ Appendix C Tensor networks
In order to compare between the Wilson and staggered fermion discretization approaches, we carry out a continuum limit extrapolation. This implies following lines of constant physics in the physical phase space, which requires the calculation of the mass shift for each $N,x,l_{0}$. With our work confined to a fixed volume of $N/\sqrt{x}=30$, this necessitates the determination of the mass shift in Eq. (5) for individual values of $N$ and $l_{0}$ only. In this work, we followd the method based on the electric field density introduced in Ref. [26]. Table 1 complements Fig. 6 and quantifies the effect of not accounting for an $l_{0}$ dependence on the mass shift with staggered fermions. Finally, Fig. 10 is qualitatively showing the effect of not accounting for the mass shift when using Wilson fermions, which was also observed in [26]. The bond dimension for all MPS data was set to $D=20,40,60,80$ and we then extrapolate the last 3 points to infinite $D$ in the electric field density against $1/D$ plot as shown in Fig. 8. The central value is taken to be the average of the extrapolated value of the linear fit and the electric field density value at the highest $D$ calculated, while the error is approximated to be half the difference between these values. The smaller $D$ solutions were given as initial ansatz to the higher $D$, while $D=20$ was initialized with a random MPS.
@@
Figure 8: Electric field density $\langle L_{W}\rangle$ against inverse bond dimension $1/D$ to extrapolate to $D\to\infty$ with $D=40,60,80$. This example is for Wilson fermions at $N=100$, $l_{0}=0.1$, $m_{\text{lat}}/g=-0.08236266$. Similar behavior was observed for staggered fermions as well.
Figure 9: Electric field density $\langle L_{W,S}\rangle$ against $ag$ to extrapolate to the continuum limit with fixed volume $N/\sqrt{x}=30$, fixed $m_{\text{r}}/g=0.01$, using $N=70,80,90,100$. The continuous orange line is a linear fit and the dashed blue line is a second order polynomial fit. $(a-c)$ is showing Wilson fermions for $l_{0}=0.1,0.4,0.6$ respectively and $(d-f)$ the same for staggered fermions with the same $x$ as Wilson fermions and the electric field density method for the mass shift [26]. Note that the error bars are much smaller than the markers and thus, are not visible.

### 11 Mar 2024 v2

@@ I Introduction
Exploring the phase diagram of a theory is central for the understanding of the fundamental laws of physics in many disciplines ranging from condensed matter to particle physics. For instance, in the case of ferromagnetic materials, critical points and phase boundaries help to understand and design magnetic materials with various applications [1, 2]. Within superconductivity, phase diagrams allow for identifying the critical temperatures and magnetic fields required for materials to exhibit zero electrical resistance [3]. In particle physics, the phase diagram of Quantum Chromodynamics (QCD) elucidates the behavior of matter under extreme conditions, relevant for the early universe and neutron stars [4]. In high-energy physics, topological terms play an important role in the study of interesting phenomena such as the breaking of charge conjugation-parity symmetry in QCD [5], and the occurrence of out-of-equilibrium dynamical effects involving axion fields [6].
In order to explore the phase structure of a given theory, one often has to resort to numerical methods. A standard tool is Monte Carlo (MC) simulation, which has been successful for probing phase diagrams [7, 8, 9, 10], including that of lattice QCD [11]. However, the sign problem poses a barrier in certain parameter regimes [12], leaving relevant questions unanswered. Specifically for lattice QCD, large baryon chemical potentials or the presence of a topological $\theta$-term would trigger this problem [13]. Thus, variations of the conventional MC approach have been proposed to tackle this obstacle [11], but so far only with limited success. Hence, there is great interest in alternative methods that bypass the issue, such as tensor networks and quantum computing [14, 15, 16, 17]. In particular, quantum computing offers a promising alternative route [15, 16, 18], with already a number of successful demonstrations [19, 20, 21, 22, 23, 24, 25, 26].
@@
Here we address these questions by studying two fermion discretizations, staggered and Wilson fermions, in order to explore any potential advantages of one type of fermion formulation over the other. We use the variational quantum eigensolver (VQE) [37] and derive a protocol for mapping this problem on a quantum circuit. The VQE, introduced as an alternative to quantum phase estimation [38], aligns with the capabilities of current and near-term quantum devices. We test our VQE using noiseless classical simulations between 6 to 12 qubits, to identify the best possible setup regarding the ansatz and gates that would capture the relevant ground states most efficiently. After the optimal ansatz-gate combination and variational parameters are found, the ground states across the phase transition are prepared on IBM’s quantum devices. We demonstrate that using state-of-the-art error mitigation techniques—zero noise extrapolation [39], readout error mitigation [40], Pauli twirling [41] and dynamical decoupling [42]—allows for obtaining precise results from the quantum measurements. To understand the minimum system sizes required to extrapolate faithfully to the continuum limit with a quantum computer, we use matrix product states (MPS). We numerically simulate intermediate system sizes and perform the continuum extrapolation, which we compare to analytical results. Our study also shows universality for the considered observables, as both discretizations lead to the same continuum values.
@@ II.2.1 Wilson fermions
$W_{W}&=x\sum_{n=0}^{N-2}\left(X_{2n+1}X_{2n+2}+Y_{2n+1}Y_{2n+2}\right)\\ &+\left(\frac{m_{\text{lat}}}{g}\sqrt{x}+x\right)\sum_{n=0}^{N-2}\left(X_{2n}X_{2n+1}+Y_{2n}Y_{2n+1}\right)\\ &+\sum_{n=0}^{N-2}\left(l_{0}+\sum_{k=0}^{n}Q_{k}\right)^{2}.\end{aligned}$ (9)
@@ III.1 Parametric ansatz circuits for VQE
Figure 2: Decomposition of a generic SO(4) gate depending on the six parameters $\theta_{1},\dots,\theta_{6}$ (a) the $R_{XX+YY}(\theta)=R_{Z_{0}}\exp(-i\theta(XX+YY)/2)R_{Z_{0}}^{\dagger}$ (b), into CNOT and Pauli rotation gates. The $R_{Z_{0}}$ rotations in the definition of $R_{XX+YY}(\theta)$ restrict the state to the real subspace. Boxes acting on a single qubit correspond to Pauli rotation gates, $R_{P}(\alpha)=\exp{-i\alpha P/2}$ ) with $P\in\{X,Y,Z\}$. Single-qubit gates where the argument is omitted refer to rotations around an angle $\pi/2$, $R_{P}(\pi/2)$. The light blue boxes represent the parameterized gates which are $R(\alpha,\beta,\gamma)=R_{X}(\gamma)R_{Z}(\beta)R_{X}(\alpha)$ in (a) and $R_{Y}(\theta)$ in (b). Panel (c) and (d) illustrate one layer of the brick and ladder ansatz, respectively, both following a non-parametric part for preparing the initial state |$\ket{\psi_{in}}$ (yellow box). The first layer in the brick ansatz has a CNOT-depth of $4$ whereas in ladder it is $22n-$ 2, where $n$ is the number of qubits, and in both cases it increases by $4$ with each layer.
@@ IV Results
Having summarized the methods in Sec. III that we have used to explore the lattice Schwinger model, we examine the performance of the ansätze for the VQE using classical simulations. Furthermore, we demonstrate the first-order phase transition with quantum hardware using the ground states obtained from the classical simulation of the VQE. Finally, we perform the extrapolation to the continuum with data obtained from an MPS calculation and estimate the resource required for a reliable continuum limit.
@@ IV.1 VQE results
Focusing on the case of staggered fermions and using the initial states and qubit mappings discussed in Sec. III.1, the ansätze with the $R_{XX+YY}$ gates generally returned higher fidelities compared to the ones with the $SO(4)$ gates. In the latter case, we observed the tendency that the final parameters obtained by the optimization routine at the end of the VQE were all close to 00, resulting in a state formed only from the initial non-parametric part (which prepares the |$\ket{10}^{\otimes N/2}$ state). In contrast, with $R_{XX+YY}$ gates both of the architectures performed comparably well and were able to achieve fidelities above $0.99$. Since the brick architecture has smaller CNOT-depth, we again focus on this one for the staggered discretization. Considering $m_{\text{lat}}/g=10$ first, the staggered discretization behaves qualitatively similar to the Wilson fermions, as Fig. 3 reveals. We again observe the characteristic discontinuities in the electric field density (second row of Fig. 3) and the particle number (fourth row of Fig. 3), indicating the first-order quantum phase transition. Comparing to the Wilson fermions, the transition for the staggered discretization occurs at larger $l_{0}$, despite the fact that for both discretizations we match the qubit number, which effectively means that for the staggered fermions we have a larger number of physical sites and, since the volume is fixed to $N/\sqrt{x}=30$, a finer lattice spacing than for the Wilson fermions. This effect might be caused by a larger additive mass renormalization compared to the Wilson case. Also for the staggered discretization, we see the same trend that with increasing system size it is shifted to smaller values of the $\theta$-angle (for details, see Appendix D). In particular, the data from exact diagonalization is in excellent agreement with the one from simulated VQE with one ansatz layer, thus showing that our ansatz circuit together with the qubit mapping described in Sec. III is appropriate for capturing the physics of the model in the regime of large masses.
@@ IV.2 Inference runs on quantum hardware
Figure 3: Electric field density $\langle L_{W,S}\rangle$ and particle number $\langle P_{W,S}\rangle$ against $l_{0}$ with data from quantum hardware (black crosses), as compared to the noiseless expectation values (red pluses) and exact diagonalization (blue circles) for Wilson and staggered fermions respectively. The staggered fermions in this case have the same number of qubits as the Wilson fermions. Hence, while the title of each column specifies the $N$ for Wilson fermions, for staggered it is taken to be double that value. The lattice mass for these data is set to $m_{\text{lat}}/g=10$, so that we are above the second order phase transition of Fig. 1, without having to account for the mass shift. Therefore, we can observe the first-order phase transition. Note that the error bars, which are discussed in Appendix B, are much smaller than the y-scale and thus, are not visible.
@@ IV.3 Continuum extrapolations with MPS
Figure 6: Electric field density $\mathcal{F}/{g}$ against $l_{0}$ for Wilson fermions and staggered fermions at $m_{\text{r}}/g=0.01$. The following MSE values are the mean squared error of each of the data as compared to the continuum mass perturbation theory prediction for the electric field density $\mathcal{F}/g$ given in Eq. (20). For staggered fermions with $\text{MS}_{t}$ and same $x$ (number of qubits as Wilson), the MSE is $1.041\times 10^{-6}$ ($6.16\times 10^{-7}$), and respectively for $\text{MS}_{L}$ $6.143\times 10^{-9}$ ($2.617\times 10^{-8}$). For Wilson fermions the MSE is $7.926\times 10^{-9}$. The error bars emanate from the errors in the variational algorithm to compute the relevant ground states, the extrapolation in bond dimension and in lattice spacing.
@@ V Conclusion
Given that for the staggered case our MPS simulations showed better convergence towards the continuum limit, and our ansatz circuit for the VQE is slightly simpler, this suggests that the staggered discretization might be easier to address on current and near-term quantum hardware. Combining our results with some of the techniques of Ref. [19], it seems possible to reach larger lattice sizes on the order of 100 qubits, and for the future we plan to study the continuum extrapolation directly on quantum hardware. Moreover, it is an interesting question how both fermion discretizations perform in higher dimensions. While for the Schwinger model we integrated the gauge fields out, this is no longer possible for two and more spatial dimensions. This might change the performance of both discretizations and the complexity of a VQE ansatz required to reliably capture the ground state.
@@ Appendix A Derivation of the Hamiltonian
In this appendix we derive the Hamiltonian for Wilson fermions, in particular, we provide the details of the fermion ordering chosen. Subsequently, we show the Jordan Wigner transformation and the final Hamiltonians in spin formulation used in the simulations for both Wilson and staggered fermions.
@@
$+\left(\frac{m_{\text{lat}}}{g}\sqrt{x}+xr\right)\sum_{n=0}^{N-1}\left(X_{2n}X_{2n+1}+Y_{2n}Y_{2n+1}\right)$
@@ Appendix B Zero noise extrapolations
Figure 7: Electric field density $\langle L_{W,S}\rangle$ and particle number $\langle P_{W,S}\rangle$ against noise factor for zero noise extrapolations (ZNE) on the data (black crosses) obtained from quantum hardware for different parameters specified in the titles of each subplot with Wilson and staggered fermions. The red pluses are noiseless values of the observables and the blue circles are the exact diagonalization results. For the fits we fit a first-order polynomial to either all points (blue continuous line) or to points with noise factor 1, 3 (dashed orange line). We also fit a second order polynomial to all points (green dotted dashed line). The brown triangle is the weighted average of the extrapolated points of these 3 fits. The plots on the left $(a-d)$ are showing examples where the ZNE helped take the expectation values measured closer to the noiseless result, while on the right $(e-h)$ we show occasions where the ZNE might not be able to improve the result. However, overall Fig. 5 indicates that ZNE can significantly improve the results. The calculation of the error bars is described within this Appendix and where they are not visible, they are smaller than the markers and y-scale.
@@ Appendix C Tensor networks
In order to compare between the Wilson and staggered fermion discretization approaches, we carry out a continuum limit extrapolation. This implies following lines of constant physics in the physical phase space, which requires the calculation of the mass shift for each $N,x,l_{0}$. With our work confined to a fixed volume of $N/\sqrt{x}=30$, this necessitates the determination of the mass shift in Eq. (5) for individual values of $N$ and $l_{0}$ only. In this work, we followed the method based on the electric field density introduced in Ref. [28]. Table 1 complements Fig. 6 and quantifies the effect of not accounting for an $l_{0}$ dependence on the mass shift with staggered fermions. Finally, Fig. 10 is qualitatively showing the effect of not accounting for the mass shift when using Wilson fermions, which was also observed in [28]. The bond dimension for all MPS data was set to $D=20,40,60,80$ and we then extrapolate the last 3 points to infinite $D$ in the electric field density against $1/D$ plot as shown in Fig. 8. The central value is taken to be the average of the extrapolated value of the linear fit and the electric field density value at the highest $D$ calculated, while the error is approximated to be half the difference between these values. The smaller $D$ solutions were given as initial ansatz to the higher $D$, while $D=20$ was initialized with a random MPS.
@@
Figure 8: Electric field density $\langle L_{W}\rangle$ against inverse bond dimension $1/D$ to extrapolate to $D\to\infty$ with $D=40,60,80$. This example is for Wilson fermions at $N=100$, $l_{0}=0.1$, $m_{\text{lat}}/g=-0.08236266$. Similar behavior was observed for staggered fermions as well. The error bars are emanating from the variational algorithm to compute the relevant ground states.
Figure 9: Electric field density $\langle L_{W,S}\rangle$ against $ag$ to extrapolate to the continuum limit with fixed volume $N/\sqrt{x}=30$, fixed $m_{\text{r}}/g=0.01$, using $N=70,80,90,100$. The continuous orange line is a linear fit and the dashed blue line is a second order polynomial fit. $(a-c)$ is showing Wilson fermions for $l_{0}=0.1,0.4,0.6$ respectively and $(d-f)$ the same for staggered fermions with the same $x$ as Wilson fermions and the electric field density method for the mass shift [28]. The error bars emanate from the variational algorithm to compute the relevant ground states and the extrapolation in bond dimension. They are much smaller than the markers and thus, are not visible.
Figure 10: Electric field density $\mathcal{F}/g$ against $l_{0}$ for Wilson fermions using no mass shift at $m_{\text{r}}/g=m_{\text{lat}}/g=0.01$. The MSE is calculated as in Fig. 6 and found here to be $7.939\cdot 10^{-4}$. The error bars emanate from the errors in the variational algorithm to compute the relevant ground states, the extrapolation in bond dimension and in lattice spacing.
Appendix D Lattice effects on the phase transition
[…]
$E_{l_{0}<l_{0}^{*}}=a(N-1)g^{2}\frac{l_{0}^{2}}{2},$ (27)
$E_{l_{0}>l_{0}^{*}}=a(N-1)g^{2}\frac{(l_{0}-1)^{2}}{2}+2m_{r}.$ (28)
$l_{0}^{*}=2\frac{m_{\text{r}}}{g}\frac{\sqrt{x}}{N-1}+\frac{1}{2}.$ (29)
In the limit of large $N$, Eq. (29) becomes
$l_{0}^{*}\approx 2\frac{m_{\text{r}}}{g}\frac{\sqrt{x}}{N}+\frac{1}{2},$ (30)
The mass shift decreases with decreasing $ag$ [28, 53]. Further, the factor $1/(1-ag)$ decreases as $ag\to 0$, which leads to an overall decrease of the first term and Eq. (31) tends towards the continuum limit. Hence, we see that the transition point on finite lattices $l_{0}^{*}$ is generally shifted towards values larger than $1/2$ and gets closer to the continuum theory prediction for decreasing lattice spacing, as observed in Sec. IV.2 and supported by Fig. 3.